# Bifurcations in the Kuramoto model with external forcing and higher-order interactions

### v1

@@ Abstract
Synchronization is an important phenomenon in a wide variety of systems comprising interacting oscillatory units, whether natural (like neurons, biochemical reactions, cardiac cells) or artificial (like metronomes, power grids, Josephson junctions). The Kuramoto model provides a simple description of these systems and has been useful in their mathematical exploration. Here we investigate this model in the presence of two characteristics that may be important in applications: an external periodic influence and higher-order interactions among the units. The combination of these ingredients leads to a very rich bifurcation scenario in the dynamics of the order parameter that describes phase transitions. Our theoretical calculations are validated by numerical simulations.
@@ I Introduction
In a seminal paper published in 1975, Kuramoto showed that non-linear oscillators with a stable limit cycle could synchronize their motion if the mutual interactions were sufficiently weak [1]. In his model, the oscillators were characterized by a single variable, describing the phase along the cycle, and pairwise interactions were approximated by smooth functions of the phase difference between the oscillators [2, 3, 4]. Kuramoto showed that synchronization would occur only if the coupling strength was larger than a minimum value, and would increase smoothly for larger coupling intensities, as in a continuous phase transition.
The Kuramoto model has become a paradigm in the field of synchronization and has been explored and extended in many directions since its inception [5]. Examples are the introduction of frustration [6], networks of connections and different distributions of the oscillator’s natural frequencies [7, 8, 9] and multi-dimensional models [10, 11, 12, 13]. One particularly important, and largely overlooked extension, considers the dynamics of periodically forced oscillators, where mutual synchronization competes with forced entrainment [14, 15, 16]. In many biological and artificial systems synchronization can indeed be strongly affected by external stimuli. It is well known that photo-sensitivity to flashing lights can trigger epileptic seizures [17] and that artificial pacemakers can control synchronization of heart cells [18]. Information processing in the brain can also be triggered by visual, auditory or olfactory inputs [19]. The visual cortex, in particular, exhibits different patterns of synchronized neuronal firing when subjected to stimuli [20].
This model considers $N$ oscillators described by their phases and interacting with each other according to
@@
where $K$ is a coupling constant and the natural frequencies $\omega_{i}$ are drawn at random from a distribution $g(\omega)$ of mean $\omega_{0}$ and unit variance (this can always be achieved by a proper scaling of the variables). It has been shown that a phase transition exists in the large-$N$ limit, with oscillators becoming synchronized as soon as $K$ is larger than some critical value $K_{c}$ that depends on $g(\omega)$. The transition may be of first or of second order, again depending on $g(\omega)$.
@@
The introduction of an external forcing, acting on each oscillator as $F\sin(\sigma t-\theta_{i})$, is an important extension of the original Kuramoto model [21, 22, 14, 23]. This corresponds to an overall influence of the environment upon the system. It creates a tendency to synchronize at the frequency $\sigma$, and therefore competes with the internal dynamics if $\sigma\neq\omega_{0}$. This problem was studied by Childs and Strogatz [14], who found that the asymptotic state can undergo several phase transitions as the parameters $F$ and $\Omega=\sigma-\omega_{0}$ are varied. Concretely, they have found explicit formulas for the boundaries between different phases in the $(F,\Omega)$ plane. These boundaries are bifurcations curves, which may be of Hopf, saddle-node, SNIPER (saddle-node with infinite period) or homoclinic type.
@@
More recently, attention has shifted to the effects of higher order interactions in systems of coupled oscillators [24]. In many complex systems, interactions go beyond pairwise terms, and involve the simultaneous action of groups of agents. Examples can be found in neuroscience [25, 26, 27, 28, 29], ecology [30, 31], biology [32] and social sciences [33, 34]. Studies have shown that the inclusion of higher order interactions have important consequences in the propagation of epidemics [35, 36, 37] and synchronization [38, 39, 40, 41, 42, 43, 44, 45].
@@
Pairwise, or two-body, interactions corresponds to the notion of an “interaction graph”. For higher order interactions this concept is replaced by that of a simplicial complex or hypergraphs which, besides nodes and edges, also contains triangles, tetrahedra, etc. This is usually done by the addition, in the Kuramoto equation for $\dot{\theta}_{i}$, of interaction terms proportional to $K_{2}\sin{(2\theta_{j}-\theta_{k}-\theta_{i})}$ and $K_{3}\sin{(\theta_{j}+\theta_{k}-\theta_{m}-\theta_{i})}$. In such a way each oscillator is allowed to interact with two or three of its neighbours simultaneously, in a way which is not reducible to two-body terms. These particular types of interactions were investigated by Skardal and Arenas [40], who showed that, if the higher-order interactions are strong enough, they can lead to bi-stability and the appearance of a first-order phase transition. The corresponding bifurcation diagram, which contains a saddle-node line and a pitchfork line, was also discussed in this work.
@@
Here we bring together the two generalizations discussed above, treating a Kuramoto system which is subject to an external influence and whose internal dynamics contains higher-order terms of the form considered in [40]. It turns out that, as already observed in previous works, the dynamics of the system depends on $K_{2}$ and $K_{3}$ only through their sum, $K_{23}=K_{2}+K_{3}$. Exploring the dynamics in full detail on the four-dimensional space of parameters $(K_{1},K_{23},F,\Omega)$ is impractical, so we develop two approaches: we investigate the effect of $K_{23}$ on the bifurcation diagram of the forced system, and the effect of the forcing on the bifurcation diagram of the simplicial system.
We find that any positive amount of forcing changes the nature of the pitchfork curve on the $(K_{1},K_{23})$ plane, turning it into a saddle-node curve. Moreover, the bifurcation curves are changed in such a way that with increasing forcing the bi-stability region shrinks and eventually disappears for a critical value of $F$. On the other hand, the influence of $K_{23}$ on the bifurcation curves on the $(F,\Omega)$ plane is even more dramatic. In particular, all bifurcation curves are doubled if $K_{1}<2$ and $K_{23}$ is larger than some critical value.
@@ II The model
$+$ $\frac{k_{3}}{N^{3}}\sum_{j,k,m=1}^{N}\sin{(\theta_{j}+\theta_{k}-\theta_{m}-\theta_{i})}+F\sin(\sigma t-\theta_{i}),$
where $K_{1}$ is the usual coupling constant, while $k_{2}$ and $k_{3}$ are new coupling constants that control the relative intensity of $2$-simplexes and $3$-simplexes, respectively. The natural frequencies $\omega_{i}$ are still drawn at random from a distribution $g(\omega)$ of mean $\omega_{0}$ and unit variance. The particular form of higher order interactions described above is the same as in [40] and we refer to it as asymmetric because of the way $\theta_{i}$ enters the equations. Symmetric higher order interactions have also been considered [39, 41] but we shall not discuss them here.
@@
$z=re^{i\psi}\equiv\frac{1}{N}\sum_{i=1}^{N}e^{i\theta_{i}},$ (3)
@@
$\dot{\theta}_{i}=(\omega_{i}-\sigma)+K_{1}r\sin(\psi-\theta_{i})+k_{2}rr_{2}\sin{(\psi_{2}-\theta_{i}-\psi)}+k_{3}r^{3}\sin(\psi-\theta_{i})-F\sin(\theta_{i}).$ (5)
@@
Upon introducing the well known Ott-Antonsen ansatz,
@@
this system collapses into two equations, namely
@@ III.1 Saddle-Node bifurcations
Saddle-node (SN) bifurcations are characterized by one of the eigenvalues of the Jacobian being zero at the equilibrium solutions $(r_{0},\psi_{0})$. The bifurcation manifold is thus determined by the equations $a(r_{0},\psi_{0})=0$, $b(r_{0},\psi_{0})=0$, and $\det[J(r_{0},\psi_{0})]=0.$ Solving these for $F$, $\Omega$ and $\psi$ leads to parametric equations in terms of the order parameter $r$:
@@ III.2 Hopf bifurcations
III.3 Example: $K=1$ and $K_{23}=7$
@@
Figure 1 shows the saddle-node and Hopf curves for $K=1$ and $K_{23}=7$. A dramatic difference with the case $K_{23}=0$ is that there are now two bifurcation “tongues” where there was previously only one: we find two saddle-node curves, two SNIPER curves and two Hopf curves. When $K_{23}=0$ (only pairwise interactions) all these bifurcation manifolds exist only for $K_{1}>2$, as can be seen from Eq.(25). That means that in the absence of forcing and higher order interactions the system recovers the Kuramoto model, which synchronizes only if $K_{1}>2$ for Lorenzian distributions with $\Delta=1$. In the next section we discuss the limit of small forcing with higher order interactions. Figure 1(b) shows a heat map of the time averaged order parameter $r$, showing very clearly the phase transitions that take place as one crosses the bifurcation lines.
Figure 1: (a) Saddle-node bifurcation (thin black and thick green solid lines) and Hopf curves (red and blue dashed lines) in the $F\times\Omega$ plane for $K_{1}=1$ and $K_{23}=7$. Red points correspond to phase diagrams in figure 2. Black squares are Takens-Bogdanov bifurcation points, where the Hopf and saddle-node lines meet. (b) Heat map of the time averaged order parameter $\langle r\rangle$ showing the phase transitions promoted by the bifurcation lines.
Figure 1(a) also shows a grid with 8 selected points: four with $\Omega=0.2$ and four with $\Omega=1$. Plots of vector fields for these points in the $x=r\cos\psi$ and $y=r\sin\psi$ plane are shown in Figure 2, illustrating how limit cycles and fixed points evolve as $F$ and $\Omega$ change for fixed $K_{1}=1$ and $K_{23}=7$. Panels (a)-(d) have $\Omega=0.2$ and (e)-(h) have $\Omega=1$. Panels (d) and (h) for $F=0.02$ and $F=0.3$ respectively, show bi-stability: a vanishing order parameter, $r=0$, is a stable fixed point, but there is also a stable limit cycle with large $r$ and an unstable cycle between these two. The outer cycle represents mutual entrainment, since we are in the frame that rotates with the external force.
Increasing $F$, panel (c) for $\Omega=0.2$ and $F=0.17$, shows that the unstable limit cycle has gone through a SNIPER bifurcation whereas in panel (g), for $\Omega=1$ and $F=0.8$, the unstable limit cycle has collided with the origin in a sub-critical Hopf bifurcation, where the origin becomes unstable. Only the external limit cycle is now stable.
Increasing $F$ even more, for $\Omega=0.2$ and $F=0.3$, panel (b), two new bifurcations have occurred with respect to $F=0.17$: a SNIPER in the outer limit cycle and a saddle-node involving the origin and one of the points created at the SNIPER of the old unstable cycle. Three fixed points remain: one stable and one unstable from the SNIPER and another unstable point. For $\Omega=1$ and $F=0.99$, panel (f), we see the same structure, although the two unstable points are very close to each other and will soon disappear in another saddle-node bifurcation.
Finally, panels (a) and (e) for large $F$ shows a single stable point corresponding to forced entrainment.
Figure 2: Vector fields for the 8 red points in figure 1. Panels (a) to (d) have $\Omega=0.2$ and $F$ equal to 0.5, 0.3, 0.17 and 0.02 respectively. Panels (e) to (h) have $\Omega=1.0$ and $F$ equal to 0.3, 0.8, 0.99 and 1.2 respectively.
III.4 Evolution of Bifurcation curves

### v2

@@ Abstract
Synchronization is an important phenomenon in a wide variety of systems comprising interacting oscillatory units, whether natural (like neurons, biochemical reactions, cardiac cells) or artificial (like metronomes, power grids, Josephson junctions). The Kuramoto model provides a simple description of these systems and has been useful in their mathematical exploration. Here we investigate this model combining two common features that have been observed in many systems: external periodic forcing and higher-order interactions among the elements. We show that the combination of these ingredients leads to a very rich bifurcation scenario that produces 11 different asymptotic states of the system, with competition between forced and spontaneous synchronization. We found, in particular, that saddle-node, Hopf and homoclinic manifolds are duplicated in regions of parameter space where the unforced system displays bi-stability.
@@ I Introduction
In an influential paper published in 1967, Winfree argued that non-linear oscillators with a stable limit cycle could be characterized only by their position, or phase, along the cycle [1]. Using approximations and numerical methods, he showed that when groups of such oscillators are coupled together, the system could undergo a transition to phase synchronization, paving the way for understanding synchrony in a variety of biological systems. Later in 1975, Kuramoto expanded on Winfree’s ideas and proposed a simple model of synchronization that could be analysed in detail in the mean field limit, where the number of oscillators goes to infinity [2]. Kuramoto showed that synchronization would occur only if the coupling strength was larger than a minimum value, and would increase smoothly for larger coupling intensities, as in a continuous phase transition.
The Kuramoto model became a paradigm in the field of synchronization and has been explored and extended in many directions since its inception [3]. Examples are the introduction of frustration [4], networks of connections [5], different distributions of the oscillator’s natural frequencies [6, 7, 8] and multi-dimensional models [9, 10, 11, 12]. One particularly important, and largely overlooked extension, considers the dynamics of periodically forced oscillators, where mutual synchronization competes with forced entrainment [13, 14, 15]. In many biological and artificial systems synchronization can be strongly affected by external stimuli. It is well known that photo-sensitivity to flashing lights can trigger epileptic seizures [16] and that artificial pacemakers can control synchronization of heart cells [17]. Information processing in the brain can also be triggered by visual, auditory or olfactory inputs [18]. The visual cortex, in particular, exhibits different patterns of synchronized neuronal firing when subjected to stimuli [19].
The Kuramoto model considers a set of $N$ oscillators described only by their phases and interacting with each other according to
@@
where $K$ is a coupling constant. The natural frequencies $\omega_{i}$ are usually drawn from an unimodal symmetric distribution $g(\omega)$ centered on $\omega_{0}$, although other arbitrary choices, such as identical oscillators, bimodal or evenly spaced frequencies [20], have been considered. Kuramoto showed that a phase transition exists in the large-$N$ limit, with oscillators becoming synchronized at frequency $\omega_{0}$ as soon as $K$ is larger than some critical value $K_{c}$ that depends on $g(\omega)$. The transition may be of first or of second order, again depending on $g(\omega)$.
@@
The introduction of an external forcing, acting on each oscillator as $F\sin(\sigma t-\theta_{i})$, corresponds to an overall influence of the environment upon the system [21, 13, 22, 23, 24, 25]. It creates a tendency to synchronize at the frequency $\sigma$ and, therefore, competes with the internal dynamics if $\sigma\neq\omega_{0}$. This problem was studied by Childs and Strogatz [13], who found that the asymptotic state can undergo several phase transitions as the parameters $F$ and $\Omega=\sigma-\omega_{0}$ are varied. Concretely, they found explicit formulas for the boundaries between different phases in the $(F,\Omega)$ plane. These boundaries are bifurcations curves, which may be of Hopf, saddle-node, SNIPER (saddle-node with infinite period) or homoclinic type.
@@
More recently, attention has shifted to the effects of higher order interactions in systems of coupled oscillators [26, 27, 28, 29, 30]. In many complex systems interactions go beyond pairwise relations, and involve the collective action of groups of three or more agents that cannot be decomposed into a sum of pairs of interactions. Examples can be found in neuroscience [31, 32, 33, 34, 35], ecology [36, 37], biology [38] and social sciences [39, 40]. Studies have shown that the inclusion of higher order interactions have important consequences in the propagation of epidemics [41, 42, 43] and synchronization [44, 45, 46, 47, 48, 49, 50, 51].
@@
Pairwise, or two-body, interactions corresponds to the notion of an “interaction graph”. For higher order interactions this concept is replaced by that of a simplicial complex or hypergraph which, besides nodes and edges, also contains triangles, tetrahedra, etc. In the Kuramoto model, 3-body interactions proportional to $\sin{(2\theta_{j}-\theta_{k}-\theta_{i})}$ (asymmetric coupling) or $\sin{(\theta_{j}+\theta_{k}-2\theta_{i})}$ (symmetric coupling), for example, can be added to the equation of motion of $\dot{\theta}_{i}$ [28], and justified by phase reduction methods [52, 53]. Specific choices have been considered separately in order to simplify the analysis. Symmetric terms, for instance, have been studied in [26, 45, 47, 30] whereas asymmetric interactions were considered in [46, 29, 54, 55]. Skardal and Arenas, in particular, have considered 3 and 4-body asymmetric interactions, with coupling constants $K_{2}$ and $K_{3}$, respectively, acting along with pairwise interactions with coupling intensity $K_{1}$. They showed that if the $K_{2}$ or $K_{3}$ are large enough, they can lead to bi-stability, promoted by a saddle-node bifurcation, and the appearance of a first-order phase transition [46]. Moreover, the dynamics of the system depends on $K_{2}$ and $K_{3}$ only through their sum, $K_{23}=K_{2}+K_{3}$.
@@
Here we bring together the two features discussed above, treating a Kuramoto system in which the oscillators are subjected to an external periodic forcing as in [13] and whose internal dynamics contains higher-order terms of the form considered in [46]. Exploring the dynamics in full detail on the four-dimensional space of parameters $(K_{1},K_{23},F,\Omega)$ is impractical, so we develop two approaches: we investigate the effect of $K_{23}$ on the bifurcation diagram of the forced system using [13] as a basis, and the effect of the forcing on the bifurcation diagram of the simplicial system, comparing the results with those in [46].
We find that $K_{23}$ produces important qualitative changes on the bifurcation diagram on the $(F,\Omega)$ plane. In particular, if $K_{1}<2$ and $K_{23}$ is greater than some critical value, corresponding to regions of bi-stability in the unforced system, all bifurcation curves are doubled, partitioning the parameter space into several regions of different asymptotic behaviors. We also found that any positive amount of forcing changes the nature of the pitchfork curve on the $(K_{1},K_{23})$ plane, turning it into saddle-node curves. In this plane it is also possible to observe the joining of the duplicated Hopf bifurcation curves, giving rise to a Bautin critical point, where a line of saddle-node bifurcation of cycles starts.
The rest of the paper is organized as follows: in Section II, we describe the Kuramoto model with external forcing and higher-order interactions, deriving the dynamical equations for the order parameter using the Ott-Antonsen ansatz. In Section III, we calculate the bifurcations curves and describe the different asymptotic behaviors of system considering the simplicial parameters $K_{1}$ and $K_{23}$ as fixed while in Section IV we fixed the forcing parameters. Finally, in Section V we present some discussions and draw our final remarks.
@@ II The model
$+$ $\frac{K_{3}}{N^{3}}\sum_{j,k,m=1}^{N}\sin{(\theta_{j}-\theta_{k}+\theta_{m}-\theta_{i})}+F\sin(\sigma t-\theta_{i}),$
where $K_{1}$ is the usual coupling constant, while $K_{2}$ and $K_{3}$ are new coupling constants that control the relative intensity of $2$-simplexes and $3$-simplexes, respectively. The natural frequencies $\omega_{i}$ are drawn at random from a distribution $g(\omega)$ of mean $\omega_{0}$ and unit variance. The particular form of higher order interactions described above is the same as in [46] and we refer to it as asymmetric because $\theta_{j}$ and $\theta_{k}$ have distinct coefficients. Symmetric higher order interactions have also been considered [45, 47] but we shall not discuss them here.
@@
$z=re^{i\psi}\equiv\frac{1}{N}\sum_{j=1}^{N}e^{i\theta_{j}},$ (3)
@@
$\dot{\theta}_{i}=(\omega_{i}-\sigma)+K_{1}r\sin(\psi-\theta_{i})+K_{2}rr_{2}\sin{(\psi_{2}-\theta_{i}-\psi)}+K_{3}r^{3}\sin(\psi-\theta_{i})-F\sin\theta_{i}.$ (5)
@@
Upon introducing the well known Ott-Antonsen ansatz [56],
@@
this system collapses into a single equation for the complex parameter $\alpha$, namely
@@
Because changing $\Omega\to-\Omega$ is equivalent to changing $\psi\to-\psi$, we will restrict our attention to the case $\Omega>0$.
@@ III.1 Saddle-Node bifurcations
Saddle-node (SN) bifurcations are characterized by one of the eigenvalues of the Jacobian being zero at the equilibrium solutions $(r_{0},\psi_{0})$. The stability of the fixed points depends on the other eigenvalue of the Jacobian matrix: if positive they correspond to a saddle and an unstable node and, if negative, to a saddle and a stable node.
The bifurcation manifold is determined by the equations $a(r_{0},\psi_{0})=0$, $b(r_{0},\psi_{0})=0$, and $\det[J(r_{0},\psi_{0})]=0.$ Solving these for $F$, $\Omega$ and $\psi$ leads to parametric equations in terms of the order parameter $r$:
@@
A particular case of saddle-node bifurcation occurs when the fixed points appear on a limit cycle. Right before the bifurcation the limit cycle has a very large period, that becomes infinite at the bifurcation, when one of the eigenvalues of the Jacobian is zero. The bifurcation is dubbed SNIPER, for saddle-node infinite period, or SNIC, which stands for saddle-node on an invariant cycle.
@@ III.2 Hopf bifurcations
III.3 Homoclinic bifurcations
From the TB point there also emerges the homoclinic bifurcation manifold, in which a periodic orbit collides with a saddle fixed point [57]. This curve starts tangentially to the SN curve at the TB point, where both eigenvalues of the Jacobian are zero, and also ends tangentially an another special point, the saddle-node-loop point, where the SN and SNIPER curves meet. There is no theoretical framework for analytical analysis of homoclinic curves, so we performed numerical calculations using the Julia package BifurcationKit [58].
III.4 Example: $K_{1}=1$ and $K_{23}=7$
@@
Figure 1 shows the bifurcation curves for $K_{1}=1$ and $K_{23}=7$. A dramatic difference with the case $K_{23}=0$ is that there are now two bifurcation “tongues” where there was previously only one: we found two SN curves, two SNIPER curves, two Hopf curves and two homoclinic bifurcation curves (see caption for details). When $K_{23}=0$ (only pairwise interactions) all these bifurcation manifolds exist only for $K_{1}>2$, as can be seen from Eq.(25). For $K_{1}<2$ there is a single stable fixed point corresponding to forced entrainment. In the absence of forcing and higher order interactions the system recovers the Kuramoto model, which synchronizes only if $K_{1}>2$ for Lorenzian distributions with $\Delta=1$. In the next section we discuss the limit of small forcing with higher order interactions.
Figure 1: Bifurcation curves in the $(F,\Omega)$ plane for $K_{1}=1$ and $K_{23}=7$, dividing the plane into ten regions, indicated by capital letters. (a) Bifurcation diagram overview: full lines are Saddle-Node bifurcations (thin black and thick green), dotted lines are SNIPERs and dashed lines are the Hopf curves (red for super-critical and blue for sub-critical). Black squares represent Takens-Bogdanov points, where the Hopf and SN bifurcations meet. (b) Zoom of the thick green SN curve, showing in detail regions E and F. (c) Zoom on the lower Takens-Bogdanov point, with a clear view of the lower homoclinic bifurcation curve (dot-dashed pink line). (d) Zoom on the upper Takens-Bogdanov point, showing the upper homoclinic bifurcation (dot-dashed pink line). The stars mark saddle-node-loop points, where the SNIPER bifurcations turn into SN and the homoclinic bifurcation ends.
The $(F,\Omega)$ plane is divided into 11 distinct regions, indicated in Figures  1 and  3 by capital letters. In each region the oscillators can be mutually synchronized or synchronized with the external force, exhibiting single or bi-stability, with continuous or discontinuous transitions between such states. Plots of vector fields in the $x=r\cos\psi$ and $y=r\sin\psi$ plane are shown in Figure 2, illustrating the behavior of the system in each region of the phase diagram and showing how limit cycles and fixed points bifurcate as $F$ and $\Omega$ change for fixed $K_{1}=1$ and $K_{23}=7$. Panels (a)-(j) show the vector fields in the respective regions A to J. Below we provide a brief description of the asymptotic dynamics in each of these regions. We recall that the coordinate system is rotating with the frequency of the external force, meaning that fixed points represent forced entrainment.
Region A - The system exhibits bi-stability, with a stable fixed point close to $r=0$ (disordered motion) and a stable limit cycle with large amplitude, representing mutual entrainment of the oscillators. An unstable limit cycle separates these two stable solutions, as shown if Fig. 2(a).
Region B - Crossing from A to B at constant $\Omega$ passes through a sub-critical Hopf bifurcation (blue dashed line) where the unstable cycle collides with the stable fixed point, that moved from the origin, and becomes unstable. The bifurcation represents a discontinuous jump from partial synchronization with the external force to highly synchronized mutual entrainment, which is now the only stable cycle, as can be seen in Fig. 2(b).
Region C - There is only one stable fixed point, representing forced entrainment of the oscillators. It comes about from the SNIPER bifurcation (black dotted line in Fig. 1) occurring on the previously stable limit cycle that existed in region B . This is illustrated in Fig. 2(c), where the ‘ghost’ of the limit cycle is still visible.
Region D - This is the region that dominates the plane above the diagonal, where $F$ is larger than $\Omega$. The system exhibits a single stable fixed point corresponding to forced entrainment, as shown in Fig. 2(d). Moving from C to D, the unstable and saddle points shown nearby in Fig. 2(c) collide in a SN bifurcation (thin black line in Fig. 1). From B to D the stable cycle collapses with the unstable fixed point, in a (reversed) super-critical Hopf bifurcation.
Region E - Crossing from C to E involves the SN bifurcation represented by the thick green line in Fig. 1. The bifurcation creates a pair of fixed points, a saddle and an unstable one, which can be seen close to the origin in Fig. 2(e). The system exhibits bi-stability again, with a fixed point near $r=0$ (disordered motion) and a stable fixed point with large $r$ (forced entrainment).
Region F - Going from A to F or from E to F requires crossing a SNIPER bifurcation. From A to F the SNIPER happens in the unstable limit cycle and from E to F we see the collision of two points creating the stable cycle in F (reversed SNIPER). Region F is again a zone of bi-stability where disordered motion co-exists with mutual entrainment, as can be seen in Fig. 2(f).
Region G - Between F and G we cross a homoclinic bifurcation: starting from G, the unstable limit cycle collides with the saddle point and disappears. The system is bi-stable with a large amplitude limit cycle and a fixed point with $r\approx 0$, see Fig. 2(g). Region G can also be accessed from A, where a SN bifurcation creates the unstable and saddle point seen close to the unstable cycle.
Region H - Similarly to region B, the only stable state is the limit cycle, corresponding to mutual entrainment. An extra pair of unstable and saddle fixed points can be noticed in Fig. 2(h), due to the extra SN bifurcation at the boundary with region B.
Region I - From C to I we cross again a homoclinic bifurcation and a stable cycle appears close to the saddle point, as seen in Fig. 2(i). The system has bi-stability with the stable cycle, representing an oscillatory forced entrainment, plus a fixed point representing forced entrainment with constant order parameter.
Region J - The subcritical Hopf bifurcation at the boundary with I makes the cycle collide with the central unstable point, leaving an unstable fixed point behind, see Fig. 2(j). Only the fixed point representing forced entrainment with constant order parameter remains.
Figure 2: Vector fields, fixed points and periodic orbits of regions A-J shown in Figures 1-(a) to (d) for $K_{1}=1$ and $K_{23}=7$. Full (empty) circles are stable (unstable) nodes and spirals, and crosses are saddle points. Full and dashed lines are stable and unstable orbits, respectively. Specific parameter values for each panel are: (a) $F=0.3$, $\Omega=1.0$; (b) $F=0.3$, $\Omega=0.2$; (c) $F=0.8$, $\Omega=1.0$; (d) $F=1.2$, $\Omega=1.0$; (e) $F=0.17$, $\Omega=0.2$; (f) $F=0.15$, $\Omega=0.1$; (g) $F=0.2612$, $\Omega=0.347$; (h) $F=0.262$, $\Omega=0.3475$; (i) $F=1.4575$, $\Omega=1.495$; (j) $F=1.4555$, $\Omega=1.4895$.
III.5 Evolution of Bifurcation curves